OpenStudy (anonymous):

Is square root 1 minus cosine squared theta = −sin Θ true? If so, in which quadrants does angle Θ terminate?

OpenStudy (anonymous):

@dunnitagainn

OpenStudy (dunnitagainn):

Okay so this is going to be a trigonometric identity, have you covered those yet?

OpenStudy (anonymous):

Sorry, doesn't ring a bell. :/

OpenStudy (dunnitagainn):

so these can be considered the laws, and they are also good to memorize, there are a few though

OpenStudy (dunnitagainn):

just give me one second

OpenStudy (anonymous):

Okay

OpenStudy (dunnitagainn):

is the cosine^2 theta under the square root with the 1?

OpenStudy (anonymous):

yeah like this: \[\sqrt{1-\cos^2\theta}\]

OpenStudy (dunnitagainn):

okay this one is a little tricky I'm just trying to find the right way to explain it.

OpenStudy (anonymous):

Take your time :)

OpenStudy (dunnitagainn):

okay, so negative sine is the reciprocal of sine meaning cosecant. CSC theta is just 1 over sine theta. have you covered these yet? CSC, SEC, COT?

OpenStudy (anonymous):

I've seen them very little... So, what your saying is? \[\cos^{-1} \theta =\frac{ 1 }{ \sin \theta }\]

OpenStudy (dunnitagainn):

no, -cos theta and cos^-1 theta are different. the first is the reciprocal and the second is the inverse and they would perform different functions.

OpenStudy (dunnitagainn):

cos^-1 is actually equal to Arccos which is what you get on your calculator when you press 2nd then cosine.

OpenStudy (anonymous):

Ahh, okay. I see now.

OpenStudy (dunnitagainn):

so sin^-1 is different than csc as well, because the first would be arcsin (the inverse of sin)

OpenStudy (dunnitagainn):

in the problem did it say -sin, or sin^-1?

OpenStudy (anonymous):

it sad -sin

OpenStudy (dunnitagainn):

okay, so sqrt(1-cos^2theta) = -sin theta we know because of a pythagorean identity that sin^2theta+cos^2theta=1, we can rewrite that to say cos^2theta= 1-sin^2theta because we want to use the same parts here, so we want all sin so now we have: sqrt(1-1-sin^2theta)= -(sin) theta we are going to square both sides to get rid of the square root, now we have 1-1-sin^2theta = -sin^2theta the ones cancel out so we are left with -sin^2theta = -sin^theta so we know they are equal

OpenStudy (dunnitagainn):

that last part is supposed to be -sin^2theta

OpenStudy (dunnitagainn):

and then since we know that sin is negative and sin= y, and y is negative in the third and fourth quadrants, we know that it must terminate in one of the two. and since earlier cosine was positive, i think we can assume the fourth, i am unsure about that part though

OpenStudy (dunnitagainn):

here is a link that should be really helpful http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf

OpenStudy (dunnitagainn):

did that all make sense??